OpenStudy (anonymous):

Please help i dont understand how to do this For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. 0.08 0.496 2.48 5

OpenStudy (prettygirl_shynice):

whar u think it is

OpenStudy (anonymous):

i dont know

OpenStudy (welshfella):

find the value of f(x) when x = 3 and then when x = 8 these are labelled f(3) and f(8) can you do that?

OpenStudy (anonymous):

so its f3=0.01(2)3?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

though write it f(3)

OpenStudy (anonymous):

and i multiply it all together? i dont remember how to do this

OpenStudy (welshfella):

hold on I think the function is (2)^x

OpenStudy (anonymous):

it is my bad

OpenStudy (welshfella):

so f(3) = 0.01 * 2^3

OpenStudy (anonymous):

yes sorry

OpenStudy (welshfella):

that = 0.01 * 2 * 2 * 2

OpenStudy (welshfella):

the exponent 3 means * 2 3 times

OpenStudy (anonymous):

.08?

OpenStudy (anonymous):

0.08

OpenStudy (welshfella):

right now do the same with x = 8

OpenStudy (welshfella):

use ur calulator for this one

OpenStudy (anonymous):

f(8)=0.01(2) ^8

OpenStudy (welshfella):

thats right

OpenStudy (anonymous):

256?

OpenStudy (welshfella):

no 2^8 = 256 now multiply by 0.01

OpenStudy (anonymous):

thats what i meants

OpenStudy (welshfella):

??

OpenStudy (anonymous):

2.56

OpenStudy (welshfella):

256 8 0.01 = 2.56

OpenStudy (welshfella):

ok now average rate of change = change in value of f(x) / change in value of x = (2.56 - 0.08) / (8 - 3)

OpenStudy (anonymous):

2.48/5 0.496

OpenStudy (welshfella):

right

OpenStudy (anonymous):

okay thanks

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

plz press best respose if you are happy with the help

OpenStudy (anonymous):

sorry i got the question right thanks

OpenStudy (welshfella):

yw